OpenStudy (anonymous):

What is 2 to the power of 4 over 3 equal to?

OpenStudy (anonymous):

5.3

OpenStudy (anonymous):

3^x = 2 Take the natural log (ln) of both sides ln (3^x) = ln(2) Use this property of logs ln x^2 = 2 ln x x ln(3) = ln(2) Divide both sides with ln(3) x = ln(2)/ln(3) This is the exact answer x = 0.63093 x is approximately equal to 0.63093

OpenStudy (anonymous):

it needs to be square root form

OpenStudy (anonymous):

\[\sqrt[3]{8}\]